# On the Importance of Uncertainty in Decision-Making with Large Language Models

### v1

@@ 5.2 Results
Figure 2: Average regret ($\pm$ std. err.) obtained on the toxic content detection bandit task.
Figure 3: CDF of the regret obtained after observing 3200 data points on the toxic content detection bandit task (one point for each random seed).
The experimental results (20 random runs, $T=100$) are shown in Figure 3. These results empirically exhibit the importance of actively using epistemic uncertainty in contextual bandit problems with Large Language Models. Overall, we notice how the approaches leveraging epistemic uncertainty tend to achieve lower regret compared to the greedy policy. Also, we notice that the confidence interval of the greedy policy is larger compared to the ones of the TS policies. This suggests that there is a high variance in the results obtained by the greedy policy. In Figure 3, we show the empirical cumulative distribution function (CDF) of the cumulative regret. From this Figure we notice that, while there are cases where greedy performs well, there are worse cases where greedy gets a very high regret. To further investigate the cause of this finding, let us define the action selection ratio as $s_{T}(a)=\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}(a_{t}=a)$. We selected two sample runs and plotted the action selection ratio for the action “publish” during these two sample runs in Figure 4. Notice that the action “publish” is a risky action, meaning that it can lead to a strongly negative reward if toxic content is published, while the action “not publish” is safer and gives a constant reward. Figure 4 clearly shows that there are cases where the greedy policy suffers from under-exploration and persists in choosing a suboptimal arm, while Thompson Sampling (in this case, Diag. LA) exhibits a more balanced behavior. Due to this issue, the greedy algorithm will incur constant regret in the random runs which require more balanced exploration. This issue makes the greedy policy perform worse than TS policies in our experiments.
Among the Thompson Sampling techniques, the Dropout method delivers strong results even without tuning the dropout probability. In Appendix B, we investigated the effect of changing the dropout probability.
@@
Regarding the Epinet TS, we see that it has a lower average regret compared to greedy but with a larger variance compared to other TS policies. From Figure 3, we notice that this variance is due to a high worst-case regret. We conjecture that the choice of the Epinet architecture may influence results. Indeed, Epinets are a very broad class of neural networks, and exploring Epinet design was not the focus of this work. Hence, we did not perform an extensive architecture search, but we believe that our work can shed light on the potential of epinets and we leave further investigation to future work.
@@
Figure 4: Action selection ratio for the action “publish” for two particular sample runs.
Additional experiment on the IMDb dataset
In order to show how our findings generalize across datasets, we conducted an additional experiment on a different dataset (IMDb (Maas et al., 2011)), which we describe in Appendix C.
6 Prior Work
@@
The aforementioned approaches typically assume that a sufficient amount of data has been collected a priori to train an effective policy. They do not explicitly address the exploration-exploitation trade-off and they do not systematically explore issues around the role of epistemic uncertainty in decision-making. In contrast, in our work, we explicitly consider the problem of learning from interaction, and we focus on one of the most fundamental and natural decision-making tasks, which is the one of contextual bandits (Chapelle & Li, 2011). We comprehensively investigate the role of epistemic uncertainty for bandit models with pre-trained LLMs.
@@
7 Conclusion
In this paper, we investigated the role of epistemic uncertainty estimation in decision-making tasks that use natural language as input. For such tasks, using Large Language Models as agents has become the norm. However, none of the recent approaches estimates the epistemic uncertainty of the agent. We focused on a fundamental decision-making task: the contextual bandit problem, where context consists of text. We approached the bandit task with a deep regression model initialized with a pre-trained LLM. As a representative of the approaches with no uncertainty estimation, we considered an LLM bandit with a greedy policy, which picks the action corresponding to the largest predicted reward. We compared the greedy baseline with various approaches integrating uncertainty estimates into the decision process via Thompson Sampling. We adapted several epistemic uncertainty estimation techniques to LLMs, such as dropout, Laplace Approximation, and epinets. Finally, we provided an empirical analysis of bandit learning with LLMs on real-world data. Our experiments showed that using uncertainty information leads to greatly improved performance over the greedy approach. These improvements highlight the benefits of modeling uncertainty for exploration in bandit problems with text and Large Language Models. Our work suggests that uncertainty should play a more central role in developing LLM-based agents for decision-making.
Appendix A Additional details on Laplace Approximation
@@
Appendix B Additional experimental details
B.1 Experimental setting
B.1.1 Computational resources
@@
Our experiments were conducted on one NVIDIA A100 GPU with 40GBs of VRAM. One random run, with, $T=100$ took about 30 minutes for the fastest bandits (i.e., greedy and dropout) and up to 40 minutes for the bandits that had to compute the Hessian (i.e., Diag. LA and Last LA). We ran 5 different bandit models for 20 random runs during the testing phase, taking a total of around 58 hours of computation.
B.1.2 Data
@@
We use an open-source dataset, called “measuring hate speech”, which is also openly available on HuggingFace Datasets [^7] https://huggingface.co/datasets/ucberkeley-dlab/measuring-hate-speech. This dataset consists of around 136,000 comments. Each of them is associated with a continuous score, called the “hate speech score”, provided for each comment. Comments with a score $>0.5$ are considered “toxic” (around $36\%$ of the comments are labeled as toxic, while the others are not toxic). The text is tokenized with the GPT2 Tokenizer. In particular, we rely on the implementation provided by HuggingFace [^8] https://huggingface.co/docs/transformers/model_doc/gpt2#transformers.GPT2Tokenizer. We selected 128 as the maximum number of tokens for each comment (only 0.3% of the total comments were longer than 128 tokens), and we applied left padding for the comments with lengths less than 128.
B.1.3 Task
@@
We evaluate the bandit policies on an automated content moderation task. The task is framed as a contextual bandit problem where the context is the text comment, and the actions available to the agent are “not publish” or “publish”. For each time step, the agent will observe a batch of $B=32$ comments. If the agent decides not to publish a comment, a reward of 0.5 is observed regardless of the actual toxicity of the comment. If the agent publishes a non-toxic comment, a reward of 1 is observed. If the agent publishes a toxic comment, a reward of -0.5 is observed. This asymmetric reward function represents a possible example suitable for this real-world scenario. The goal of the agent is to minimize the regret (compared to the optimal publishing policy) over time by learning to make optimal publish/not publish decisions based on the text content.
B.1.4 Bandit models
@@
To investigate the role of uncertainty in LLM bandits, we compare different TS variants with the greedy baseline. Every bandit model is initialized as in Eq. 1, with a pre-trained GPT2 model with 124M parameters (Radford et al., 2019). We use the implementation provided by the HuggingFace library [^9] https://huggingface.co/docs/transformers/model_doc/gpt2. GPT2 is a Causal Language Model, which means that the attention of each token can only look to the current token and the previous ones. For our purposes, we remove the final classification head, and we take the embedded features of the last token. This is because the last token can look to all the tokens in the sentence with the attention mechanism. In the GPT2 model, this final feature vector has a dimension of 768. We then add a final linear layer with 2 output neurons, which will be trained to solve our regression task. All the parameters are updated during training. We train each model at the end of each time step for 50 epochs with the Adam optimizer (Kingma & Ba, 2014), with learning rate set to $3\cdot 10^{-5}.$ As TS variants, we include dropout, Diagonal Fisher LA (which we will call Diag. LA), LA with full Hessian on the last layer (which we will call Last LA), and Epinet TS. In the following paragraphs, we describe in further detail the different TS bandit models.
@@ B.2 Hyperparameter tuning
For each model, hyperparameters are tuned on 10 random runs (with different seeds than the testing ones) with $T=50$. The set of hyperparameters are shown in Table 1. In particular, for all the hyperparameter configurations, we selected the ones that best performed on average, measured by cumulative regret, except for the LA TS models. For the LA TS models, we selected the best configuration and the ones that are not statistically different from the best (measured with a t-test, p-value=0.05). Among those configurations, we selected the ones with the highest values for prior and observation variances, in order to induce exploration.
B.3 Additional ablation experiment on dropout
@@
Figure 5: Average regret obtained with different dropout values
@@
Among the Thompson Sampling techniques, the Dropout method delivers strong results even without tuning the dropout probability. By relying on the same dropout rate used during pre-training, we are using the same uncertainty that the original model had in learning to generate natural language. Therefore, it is not necessarily the best dropout probability to use in a bandit task. To investigate the importance of this hyperparameter in our experiments, we show dropout TS policies across various dropout probabilities in Figure 5. From these results, the rate originally used for pre-training appears optimal for the bandit task. Also, we notice that, for a smaller value of $p$ ($p=0.05$), the reduced exploration induces a larger variance, as expected.
Appendix C Additional experiment on the IMDb dataset
In this section, we describe an analogous bandit experiment we conducted on a different dataset, in order to show how our findings generalize across datasets.
C.1 Experimental setting
C.1.1 Data
We use the open-source IMDb dataset (Maas et al., 2011), which is also openly available on HuggingFace Datasets [^10] https://huggingface.co/datasets/imdb. This dataset consists of 50,000 movie reviews. Each of them is associated with a class, which can be “positive” or “negative”, according to the sentiment expressed in the review. The dataset consists of exactly $50\%$ “positive” reviews and $50\%$ “negative” reviews. The text is tokenized with the GPT2 Tokenizer. In particular, we rely on the implementation provided by HuggingFace. We selected 256 as the maximum number of tokens for each comment, and we applied left padding for the reviews with lengths less than 256.
C.1.2 Task
We evaluate the bandit policies on an automated content moderation task. The task is framed as a contextual bandit problem where the context is the text comment, and the actions available to the agent are “not publish” or “publish”. In this case, we simulate this task by considering a negative review as a toxic content, which we do not want to publish. For each time step, the agent will observe a batch of $B=32$ comments. If the agent decides not to publish a review, a reward of 0.5 is observed regardless of the actual toxicity of the content. If the agent publishes a non-toxic comment (which, in this case, is a positive review), a reward of 1 is observed. If the agent publishes a toxic comment (which, in this case, is a negative review), a reward of -0.5 is observed. This asymmetric reward function represents a possible example suitable for this real-world scenario. The goal of the agent is to minimize the regret (compared to the optimal publishing policy) over time by learning to make optimal publish/not publish decisions based on the text content.
C.1.3 Bandit models and Hyperparameter tuning
For this task, we employ the same bandit models used for the experiment shown in Section B. As for the hyperparameters, we do not tune any hyperparameter for this task specifically. Rather, we use the hyperparameters found in the hyperparameter tuning procedure for the previous experiment, shown in Section B.
C.2 Results
Figure 6: Average regret ($\pm$ std. err.) obtained on the IMDb dataset.
Figure 7: CDF of the regret obtained after observing 3200 data points on the IMDb dataset (one point for each random seed).
The experimental results (20 random runs, $T=100$) are shown in Figure 7. These results empirically exhibit the importance of actively using epistemic uncertainty in contextual bandit problems with Large Language Models also with this dataset. Overall, we notice how the greedy policy is the one that achieves the highest regret in this task. In this new task, Dropout TS and Diag. LA TS exhibit a slightly lower average regret compared to the greedy algorithm, but this improvement is within the std. err. of the greedy regret. This may be the result of the fact that we did not fine-tune the hyperparameters for this new task. Instead, we used the same hyperparameters resulted from a hyperparameter tuning on a different dataset. However, we notice that Last LA TS and Epinet TS achieve a significantly lower regret compared to the other policies. This shows that epistemic uncertainty is fundamental for a better exploration, also in this new bandit task. These findings are also confirmed by Figure 7, which shows the empirical CDF of the cumulative regrets obtained over the 20 random runs.

### v2

@@ 5.2 Results
Figure 2: Average regret ($\pm$ std. err.) obtained on the toxic bandit task.
Figure 3: Average regret ($\pm$ std. err.) obtained on the imdb bandit task.
Figure 4: Average regret ($\pm$ std. err.) obtained on the offensive bandit task.
Figure 5: Average regret ($\pm$ std. err.) obtained on the hate bandit task.
We show the experimental results (20 random runs, $T=100$) for the toxic task (Figure 3), imdb task (Figure 3), offensive task (Figure 5), and hate task (Figure 5) with GPT2 (124M) LLM bandits. These results empirically exhibit the importance of actively using epistemic uncertainty in contextual bandit problems with Large Language Models. Overall, we notice how the approaches leveraging epistemic uncertainty tend to achieve lower regret compared to the greedy policy in all the considered tasks.
Analyzing the results in more detail, we observe that Dropout TS exhibits strong performance on the toxic and hate tasks, achieving lower regret compared to the greedy policy. However, its performance is comparable to the greedy approach on the imdb and offensive tasks. This behavior is expected, as Dropout TS and the greedy policy are very similar, with the only distinction being the activation of dropout during inference in Dropout TS. It is important to notice that we did not tune the dropout probability, using the same one used during the pre-training of the base LLM. The performance of Dropout TS may possibly improve after tuning the dropout probability. In Appendix D, we investigated the effect of changing the dropout probability. Also, Dropout TS incurs no additional computational cost, requiring only a single forward pass, akin to the greedy method.
Diag. LA TS demonstrates excellent results on the toxic task and attains lower regret than the greedy policy on the offensive and hate tasks. However, its performance on the imdb task is comparable to the greedy approach. This may be attributed to the fact that hyperparameter selection was performed on the toxic task, and it may not generalize well to the imdb task.
@@
Both Last LA TS and Epinet TS consistently achieve lower regret compared to the greedy policy across all four tasks, albeit with some distinctions. Last LA TS emerges as the best-performing method in all cases, but it requires computing the full Hessian on the last layer, which can be computationally expensive. On the other hand, Epinet TS, while slightly inferior to Last LA TS, offers a more computationally efficient alternative by employing an additional small neural network. These results confirm previous findings in the literature (Osband et al., 2022; 2023b). Also, we conjecture that the choice of the epinet architecture may influence results. Indeed, epinets are a very broad class of neural networks, and exploring epinet design was not the focus of this work. Hence, we did not perform an extensive architecture search, but we believe that our work can shed light on the potential of epinets, and we will leave further investigation to future work.
@@
Figure 6: Action selection ratio for the action “publish” for two particular sample runs (toxic task).
Moreover, we notice that the confidence interval of the greedy policy is generally large. This suggests that there is a high variance in the results obtained by the greedy policy. To further investigate the cause of this finding, let us define the action selection ratio as $s_{T}(a)=\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}(a_{t}=a)$. We selected two sample runs for the toxic task and plotted the action selection ratio for the action “publish” during these two sample runs in Figure 6. Notice that the action “publish” is a risky action, meaning that it can lead to a strongly negative reward if toxic content is published, while the action “not publish” is safer and gives a constant reward. Figure 6 clearly shows that there are cases where the greedy policy suffers from under-exploration and persists in choosing a suboptimal arm, while Thompson Sampling (in this case, Diag. LA) exhibits a more balanced behavior. Due to this issue, the greedy algorithm will incur constant regret in the random runs which require more balanced exploration. This issue makes the greedy policy perform worse than TS policies in our experiments.
Experiments with 1.5B LLM
Figure 7: Average regret ($\pm$ std. err.) obtained on the hate bandit task with 1.5B LLM bandits.
To further investigate if our findings generalize to larger models, we performed additional experiments using GPT2-XL, a larger language model with 1.5B parameters. For this set of experiments, we train our models for 5 epochs for each batch of data. We perform this new set of experiments on the hate task. The results of these experiments (20 random runs, $T=100$) are shown in Figure 7. These results confirm our findings, showing how TS policies generally achieve a lower regret compared to the greedy policy. In particular, we again see how Last LA TS achieved the lowest regret among the TS variants. These results further highlight the importance of leveraging epistemic uncertainty in contextual bandit problems, even when scaling up to larger language models.
Additional ablation study on model dimension and pre-training
We also performed an additional ablation study to investigate the effect of using large pre-trained models to initialize the bandit policies in our experiments. We conduct those studies by running a bandit algorithm using a smaller transformer model with weights initialized from scratch. We report those experiments in Appendix D.4.
6 Conclusion
In this paper, we investigated the role of epistemic uncertainty estimation in decision-making tasks that use natural language as input. For such tasks, using Large Language Models as agents has become the norm. However, none of the recent approaches estimates the epistemic uncertainty of the agent. We focused on a fundamental decision-making task: the contextual bandit problem, where context consists of text. We approached the bandit task with a deep regression model initialized with a pre-trained LLM. As a representative of the approaches with no uncertainty estimation, we considered an LLM bandit with a greedy policy, which picks the action corresponding to the largest predicted reward. We compared the greedy baseline with various approaches integrating uncertainty estimates into the decision process via Thompson Sampling. We adapted several epistemic uncertainty estimation techniques to LLMs, such as dropout, Laplace Approximation, and epinets. Finally, we provided an empirical analysis of bandit learning with LLMs on real-world data. Our experiments showed that using uncertainty information leads to greatly improved performance over the greedy approach. These improvements highlight the benefits of modeling uncertainty for exploration in bandit problems with text and Large Language Models. Our work suggests that uncertainty should play a more central role in developing LLM-based agents for decision-making. As future work, a natural extension would be to use our study on LLM bandits as a stepping stone to more complex settings like text-based reinforcement learning. Another interesting direction would be to study the scaling behavior of LLM bandits in a systematic way.
Appendix A Prior Work
@@
The aforementioned approaches typically assume that a sufficient amount of data has been collected a priori to train an effective policy. They do not explicitly address the exploration-exploitation trade-off and they do not systematically explore issues around the role of epistemic uncertainty in decision-making. In contrast, in our work, we explicitly consider the problem of learning from interaction, and we focus on one of the most fundamental and natural decision-making tasks, which is the one of contextual bandits (Chapelle & Li, 2011). We comprehensively investigate the role of epistemic uncertainty for bandit models with pre-trained LLMs. The importance of exploiting epistemic uncertainty is emphasized also in (Dwaracherla et al., 2024), where they propose an approach to approximate epistemic uncertainty for the alignment problem of LLMs.
@@
Appendix B Additional details on the used loss
In this section, we elaborate in more detail the loss described in Eq. 3.
Using the loss shown in Eq. 3, we are doing MAP inference many times, applying the Bayes rule anew for each new batch of data. In this way, we perform MAP, which is a well-known established principle in deep learning, while avoiding looping throughout the whole dataset. Indeed, at a given time step $t$, the dataset will look like this: ${\mathcal{D}}=\{{\mathcal{D}}_{1},\dots,{\mathcal{D}}_{t}\}$, where each ${\mathcal{D}}_{i}$ is composed of i.i.d. data points. The posterior distribution of the parameters can be re-written as follows:
This justifies the loss in Eq. 3 as a MAP method: it is equivalent to minimizing the negative logarithm of Eq. 12.
Appendix C Additional details on Laplace Approximation
@@
Appendix D Additional experimental details
D.1 Experimental setting
D.1.1 Computational resources
@@
Our experiments were conducted on one NVIDIA A100 GPU with 80GBs of VRAM. One random run with a GPT2 (124M) bandit, $T=100$, took about 30 minutes for the fastest bandits (i.e., greedy and dropout) and up to 45 minutes for the bandits that had to compute the Hessian (i.e., Diag. LA and Last LA). We ran 5 different bandit models for 20 random runs during the testing phase, and we repeated those runs for all four tasks. In total, the experiments with GPT2 bandits took around 220 hours of computation. The experiments done with GPT2-XL (1.5B) bandits, $T=100$, 20 random runs, 5 different bandit models, hate task, took around 50 hours of computation.
D.1.2 Tasks and Data processing
We evaluate the bandit policies on various real-world tasks. These tasks consist of:
@@ D.1.2 Tasks and Data processing
  • An open-source dataset, called “measuring hate speech”, which is openly available on HuggingFace Datasets [^10] https://huggingface.co/datasets/ucberkeley-dlab/measuring-hate-speech. This dataset consists of around 136,000 comments. Each of them is associated with a continuous score, called the “hate speech score”, provided for each comment. Comments with a score $>0.5$ are considered “toxic” (around $36\%$ of the comments are labeled as toxic, while the others are not toxic). We will refer to this task as toxic.
  • An open-source dataset, called “IMDb(Maas et al., 2011), which is openly available on HuggingFace Datasets [^11] https://huggingface.co/datasets/imdb. This dataset consists of 50,000 movie reviews. Each of them is associated with a class, which can be “positive” or “negative”, according to the sentiment expressed in the review. The dataset consists of exactly $50\%$ “positive” reviews and $50\%$ “negative” reviews. We will refer to this task as imdb.
  • An open-source dataset, called “Offensive Language Identification(Zampieri et al., 2019), which is openly available on HuggingFace Datasets [^12] https://huggingface.co/datasets/tweet_eval/viewer/offensive. This dataset consists of over 14,000 English tweets. Each of them is classified as “non-offensive” or “offensive”. The dataset consists of $66.9\%$ “non-offensive” tweets and $33.1\%$ “offensive” tweets. We will refer to this task as offensive.
  • An open-source dataset, called “HatEval(Basile et al., 2019), which is openly available on HuggingFace Datasets [^13] https://huggingface.co/datasets/tweet_eval/viewer/hate. This dataset consists of 13,000 English and Spanish tweets, where some of them contain hate speech against immigrants and women. Each of them is classified as “non-hateful” or “hateful”. The dataset consists of $58\%$ “non-hateful” tweets and $42\%$ “hateful” tweets. We will refer to this task as hate.
For each task, the text is tokenized with the GPT2 Tokenizer. In particular, we rely on the implementation provided by HuggingFace [^14] https://huggingface.co/docs/transformers/model_doc/gpt2#transformers.GPT2Tokenizer. For each task, we selected a maximum number of tokens, and we applied left padding for the comments with lengths less than the selected number of tokens. For the imdb task, we selected 256 as the maximum number of tokens; for the toxic task, we selected 128 as the maximum number of tokens; for the hate and offensive tasks, we selected 64 as the maximum number of tokens. The difference is due to the difference in the length of each data point of those datasets.
@@
These tasks are framed as a contextual bandit problem as follows. The context is the text input, and the actions available to the agent are “not publish” or “publish”. For each time step, the agent will observe a batch of $B=32$ comments. The reward function is the following: if the agent decides not to publish a comment, a reward of 0.5 is observed regardless of the actual toxicity of the comment. If the agent publishes a non-toxic comment, a reward of 1 is observed. If the agent publishes a toxic comment, a reward of -0.5 is observed. This asymmetric reward function represents a possible example suitable for this real-world scenario. The goal of the agent is to minimize the regret (compared to a clairvoyant optimal publishing policy) over time by learning to make optimal publish/not publish decisions based on the text content. For the imdb task, we consider “negative” reviews as toxic content; for the offensive task, we consider “offensive” tweets as toxic content; for the hate task, we consider “hateful” tweets as toxic content.
D.1.3 Bandit models
@@
To investigate the role of uncertainty in LLM bandits, we compare different TS variants with the greedy baseline. Every bandit model is initialized as in Eq. 1, with a pre-trained GPT2 model (Radford et al., 2019) (either the base one with 124M parameters, or the XL version with 1.5B parameters). We use the implementation provided by the HuggingFace library [^15] https://huggingface.co/docs/transformers/model_doc/gpt2. GPT2 is a Causal Language Model, which means that the attention of each token can only look to the current token and the previous ones. For our purposes, we remove the final classification head, and we take the embedded features of the last token. This is because the last token can look to all the tokens in the sentence with the attention mechanism. In the GPT2 model, this final feature vector has a dimension of 768. We then add a final linear layer with 2 output neurons, which will be trained to solve our regression task. All the parameters are updated during training. We train each model at the end of each time step (for 50 epochs when we use GPT2, for 5 epochs when we use GPT2-XL) with the Adam optimizer (Kingma & Ba, 2014), with learning rate set to $3\cdot 10^{-5}.$ As TS variants, we include dropout, Diagonal Fisher LA (which we will call Diag. LA), LA with full Hessian on the last layer (which we will call Last LA), and Epinet TS. In the following paragraphs, we describe in further detail the different TS bandit models.
@@ D.2 Hyperparameter tuning
For each GPT2 (124M) bandit model, hyperparameters are tuned on 10 random runs (with different seeds than the testing ones) with $T=50$ on the toxic task. The set of hyperparameters are shown in Table 1. In particular, for all the hyperparameter configurations, we selected the ones that best performed on average, measured by cumulative regret, except for the LA TS models. For the LA TS models, we selected the best configuration and the ones that are not statistically different from the best (measured with a t-test, p-value=0.05). Among those configurations, we selected the ones with the highest values for prior and observation variances, in order to induce exploration.
D.3 Additional ablation experiment on dropout
@@
Figure 8: Average regret obtained with different dropout values.
@@
Among the Thompson Sampling techniques, the Dropout method delivers strong results even without tuning the dropout probability. By relying on the same dropout rate used during pre-training, we are using the same uncertainty that the original model had in learning to generate natural language. Therefore, it is not necessarily the best dropout probability to use in a bandit task. To investigate the importance of this hyperparameter in our experiments, we show dropout TS policies across various dropout probabilities in Figure 8 on the toxic bandit task. From these results, the rate originally used for pre-training appears optimal for the bandit task. Also, we notice that, for a smaller value of $p$ ($p=0.05$), the reduced exploration induces a larger variance, as expected.
D.4 Additional experiment on model dimension and pre-training
Figure 9: Average regret ($\pm$ std. err.) obtained on the toxic bandit task.
Figure 10: Average regret ($\pm$ std. err.) obtained on the imdb bandit task.
Figure 11: Average regret ($\pm$ std. err.) obtained on the offensive bandit task.
Figure 12: Average regret ($\pm$ std. err.) obtained on the hate bandit task.
In this experiment, we want to investigate the effect of using a large pre-trained model to initialize the bandit policies in contextual bandits with text as input. To this end, we evaluate different greedy policies on the same four tasks presented in Section 5. To investigate the role of the dimension of the model, we include a smaller transformer with a GPT2-like architecture, with 2 attention blocks, embedding dimension of 128, and 4 heads, for a total of around 7M parameters. The training procedure is analogous to the one of the pre-trained greedy policy. The results are shown in Figure 12, 12, 12, 12. These results show how using a pre-trained LLM is essential for solving contextual bandit tasks with text as input. Indeed, the policies that use a small transformer always get stuck with suboptimal actions without learning the real bandit task. By using pre-trained LLMs instead, we can leverage a model that can understand human language (since it was pre-trained on natural language generation) and focus on learning the task-specific reward function using expensive bandit data. This transfer learning approach improves the sample-efficiency of our bandit.